I would like to understand if the following equation is always, sometimes or never true and why.
(2/3)x + 4 = 2x

Answer by TimothyLamb(4379)

(2/3)x + 4 = 2x
2x - (2/3)x = 4
(4/3)x - (2/3)x = 4
(2/3)x = 4
x = 4*(3/2)
x = 6
---
x = 6 is the simplified version of the given equation, and is always true
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
Send and receive free text messages online:
https://sooeet.com